M signs a note promising to pay N Rs.3000 in 6 years with interest at 5% compounded quarterly. four years later N sells the note to P. The amount of money P pay if money worth 4% compounded semiannually is?
Answer by jorel1380(3719)

3000 x 5%,compounded quarterly x 6 years=3000 x (1+(.05/4))^(4x6)=
3000 x (1.0125)^24=Rs.4042.05
4042.05 x 4%, compounded semi-annually x 4 years=4042.05(1+(.04/2)^(2x4)=
4042.05 x (1.02)^8=Rs. 4735.91 as the final value of the note. ☺☺☺☺